Steve has a solution that is 25% antifreeze and a solution that is 65% antifreeze. How much of each should he use to obtain 12 liters of solution that is 50% antifreeze?
Answer by mananth(16949)

Steve has a solution that is 25% antifreeze and a solution that is 65% antifreeze. How much of each should he use to obtain 12 liters of solution that is 50% antifreeze?

25% 65% 50%
X 12-x 12
0.25x+0.65*(12-x)=0.5*12
0.25x+7.80 – 0.65x=6
-0.4x = -1.8
X=4.5
& 7.5 liters
mananth@hotmail.com